OpenStudy (anonymous):

F(x)=x2+8x+7 find the vertex.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

-b/2a

OpenStudy (anonymous):

what imran said. in this case b= 8, a = 1 \[\frac{-b}{2a}=-\frac{8}{2}=-4\]

OpenStudy (anonymous):

then plug in -4 for x to get \[f(-4)=(-4)^2+8\times -4=16-32+7=-16+7=-9\]

OpenStudy (anonymous):

so vertex is (-4,-9)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Thank you sO much!

OpenStudy (anonymous):

welcome!